OpenStudy (anonymous):

Three balls collide at the origin and stick together. Ball A is 4kg moving in +x direction at 3m/s towards the origin, Ball B is 6kg in + y towards origin at 5m/s and Ball C is 8kg, moving towards origin at an angle of 45 degrees from + x at 7m/s. What is the final velocity of the three balls after the collision?

OpenStudy (amistre64):

elastic versus inelastic collisions

OpenStudy (amistre64):

\[m_1v_1+m_2v_2+m_3v_3=(m_1+m_2+m_3)v\] http://www.batesville.k12.in.us/physics/apphynet/Dynamics/Collisions/inelastic_collisions.htm

OpenStudy (amistre64):



OpenStudy (amistre64):

its prolly best to use cos and sin components of each force

OpenStudy (amistre64):

bA; mv = 4(3) = 12 * cos(0) = 12 bB; mv = 6(5) = 30 * sin(90) = 30 bC; mv = -8(7) cos(45) = -28sqrt(2) -8(7) sin(45) = -28sqrt(2) 12+30-56sqrt(2) is the what we have added altogether divide this by: m1+m2+m3 to get the resulting velocity \[\frac{12+30-56\sqrt{2}}{4+5+8+8}\] since a split bC into 2 parts I assume we have to apply it as 2 parts as well

OpenStudy (amistre64):

 maybe a better representation if we go 12 over 30 up and diagonally 56 we calculate the distance from where we started .... but i cant say thats i good idea or not :)